Telescope Bootstrap Startup: 12 September 2012 UT

Observer: JHill (LBTO)
Telescope Operator: GBechetti (LBTO)
Software Support: CBiddick (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: JLittle and MMidkiff (LBTO)
Instrument Support: ESolheid (LBTO)

Summary

This night was lost to weather. There were heavy thunderstorms in the afternoon. Mountain is covered in fog.

We (software) solved the AOS communication problem in 2012G.

Rick and Olga completed MODS1 wakeup, IMCS calibration, and MODS calibration data (flats, biases and lamps).

Vidhya did a lot of testing of LBTI AO and its interactions with TCS.

Details

Preparations

The day guys found the leaking seam in the shutter door roof over the DX primary, and started work to patch it.

Testing AOS in 2012G

01:00 We switched from 2012F to a version of 2012G with AOS linked to a different boost library. That solved the AOS communication problems of Issue 4198.

We are keeping Build 2012G in place to exercise it for the rest of the night.

MODS1 IMCS calibration

02:00 RPogge on Skype and obs2 setting up for the MODS1 IMCS calibration. Olga is connected by polycom from Tucson. The MODS1 wakeup tests in the afternoon were successful. We still haven't tested the MODS1 AGw yet.

02:15 We are covered in fog at 6.4 degC. It is still mostly cloudy above the fog. Moderate wind (12 m/s) from the SW.

02:43 Moving to EL=60, LDG=207

02:51 Starting the IMCS calibration.

03:06 IMCS script completed. Rick and Olga will analyze the data tomorrow.

03:10 Test presets in TRACK and GUIDE mode

Rick's x11vnc is in ~MODSeng/Tools/bin/. Olga was using this to run obs2 from the Tucson remote room.

MODS1 Calibration Data

03:25 Olga running the calibration scripts from Tucson

Olga's calibrations continue for several hours.

LBTI AO tests

04:45 Vidhya joins us for some LBTI tests

From March 2012, M2 hexapod position for retro was X=1.72 Y=3.38 Z=4.8 RX=-30 RY=+40

We removed the +500 Tilt global offset from M3 that was added during the May 2012 observing run.

06:30 Search for source with AOIDL hexaspiral, but did not find source. Using these commands, but there are some tricks about the order in which things are compiled in IDL.
print, hexa_spiral(500e-6, 15,/apply)
from AOeng@adsecsx
aoidl
@startup-onlycom
here
.r hexapod_opt_calib
.r hexa_spiral
hexa_spiral

06:53 Why did Vidhya's command to the left M2 cause the right side preset to fail? Can we not flatten while telescope is moving? No, it was a coincidence. The error was because the WFS was powered off. See 07:51.

We found that the "On" button on the AOS GUI for the LBTI WFS actually sends a WfsOff command. See Issue 4212.

08:34 We lost all network connection to the mountain (for about 5 min). Within seconds of that, AOSL GUI put up the big red box that says the IDL license is missing. I added a note to Issue 4187

08:57:25 SX adsec goes into safe mode when we try to turn it on - this was second try.

09:00 Lost network again. Vidhya calls it a night.

More MODS1 Calibrations

09:08 Run Olga's script to take another set of calibrations while we sleep. It puts MODS to sleep when it finishes.

09:15 Humidty remains at 100% with thick fog. T=6 degC. Some rain showers are passing just south of the mountain.

-- JohnHill - 12 Sep 2012
Topic revision: r4 - 13 Sep 2012, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback